# Word Definitions from Large Language Models

### v2

@@ Abstract
1 Introduction
Many generations of readers have relied on the notion that the meaning of a word is what it says in the dictionary. Definitions of the primary and alternate senses of words have appeared in printed books well before Samuel Johnson’s famous A Dictionary of the English Language in 1755. Early western lexicography evolved from glosses and glossaries. American English was later standardized by Noah Webster in American Dictionary of the English Language in 1828.
But the primacy of dictionary definitions as the arbitrator of word meanings has come under threat by recent progress in natural language processing. Vector representations (word embeddings) have proven more valuable computationally than handcrafted semantics or definitions in describing the effective meaning of vocabulary words in a given language. And generative dialog systems like ChatGPT will happily produce full text definitions of every word when asked; indeed multiple versions of these definitions in response to different prompts.
This paper is an exploratory study of the degree of alignment between word definitions from classical dictionaries and these newer computational artifacts. There are several distinct questions we address here:
@@ 1 Introduction
  • LLMs have been described as “stochastic parrots” Bender et al. (2021). To what extent are the definitions they produce mere repetitions or even plagiarisms of classical dictionary definitions?
Our methodology is as follows. For each of over 2,500 carefully selected words, we extracted their handcrafted definitions in each of three different dictionaries (WordNet, Merriam-Webster, and Random House/Dictionary.com), and also vector representations from two prominent collections of word embeddings (FastText and Glove). We also asked two versions of ChatGPT (3.5 and 4.0) for the definitions of every word, each with two different forms of prompts. The first prompt is "what is the meaning of this word?", while the second prompt is "define this word." We measure the distances between these representations using several techniques, including vector distance/similarity, edit distance, and neighborhood correlations.
@@
  • The surface form of generated definitions differ substantially from those of published dictionaries – Although ChatGPT presumably trained in part on the handcrafted definitions in our study, little-to-no unexpected trace of these texts remain in the generated text. Indeed, we find roughly twice as many long exact matches between traditionally published dictionary definitions than we observe between published and generated texts.
@@
  • ChatGPT generated word definitions are highly accurate, consistent with published dictionaries – We evaluate SBERT definitions embeddings to identify candidates for mismatching definitions across sources. Of the 50 most distant pairs between Merriam-Webster and GPT4 definitions, human evaluation confirms that GPT4 recognized 48 compared to 34 for GPT3, and that all but one of these definitions matched the primary sense of the published dictionary.
@@
  • The consistency of LLM-generated definitions is relatively independent of frequency, unlike word embeddings – We propose a new average distance correlation metric to compare word embeddings in different dimensional spaces, and use it to study both traditional word embeddings (GloVE and FastText) and sentence embeddings of both dictionary and generated definitions. Our results suggest that SBERT embeddings of definitions (synthesized or published) may create more accurate word embeddings for low frequency words than traditional methods.
@@
This paper is organized as follows. Section 2 presents related work on dictionaries, definitions, and semantic representation. The dictionaries and generative models we use to formulate our dataset are described in Section 3. The question of just how originial generated definitions are is considered in Section 4. We use vector representations to quantify the similarity of definitions by source in Section 5, and between definitions and word embeddings in Section 6. We conclude with directions for future work in Section 7.
2 Related Work
@@
Large language models (LLM) recently burst into popular consciousness through the phenomenon of ChatGPT, which reached 100 million users in only three months Milmo (2023), offering the prospect of knowledge systems that can write professional-quality documents. Generative Pre-trained Transformer (GPT) models are a series of large language models (LLMs) developed by OpenAI. Each model in the series Radford and Narasimhan (2018) Radford et al. (2019) Brown et al. (2020) OpenAI (2023) is trained on a larger corpus of text and achieves better results on natural language processing (NLP) tasks compared to its predecessors.
ChatGPT promises to have a major impact on many fields, including healthcare Scerri and Morin and education Han et al. (2023); Mogali . The models have also been used to assess translation quality Kocmi and Federmann (2023) and the nature of human personalities Rao et al. (2023).
2.1 LLMs and Prompt Engineering
One aspect of our work is assessing how different prompts to LLMs (here, for word definitions) yield responses of different quality. Prompt engineering has been studied for Human-Computer Interaction Oppenlaender (2022), NLP Liu et al. (2021); Wang et al. (2023), and tool learning based on foundation models Qin et al. (2023). Prompt design helps overcome model bias Rao et al. (2023). Chen et al. identify the best prompt to induce correct answer from pre-trained language models at a low cost Chen et al. (2023).
Chain-of-thought prompting Wei et al. (2023) decomposes a task into a sequence of steps. It has been used to allow LLMs to use tools such as a calculator and a chemical reaction predictor Inaba et al. (2023), to improve reasoning over long input documents Sun et al. (2023), to perform better on web navigation tasks Sridhar et al. (2023), and to achieve more correct answers through dynamic sampling Aggarwal et al. (2023).
2.2 Dictionaries and Definitions
Lexicographical works can be dated back to ancient Mesopotamia Veldhuis (2019) and ancient China Creamer (1992). The ancient Greeks also use glossaries to understand religious and literature works Lee (2022); Dolezal and Velde (2021). Medieval lexicographers use derivation and alphabetization techniques Weijers (1989), which lay the foundations for early modern dictionaries Considine (2008).
The English language was standardized through dictionaries by Samuel Johnson and Noah Webster Sokolowski (2020). The Oxford English Dictionary represents another milestone, with multiple editions created in the 19th and 20th centuries. Most dictionaries in the digital age now appear in electronic form Zimmer (2014).
Researchers have shown that typical native speakers of American English know an average of 42,000 words and 4,200 multiword expressions. This number ranges from 27,000 words for the lowest 5% to 52,000 words for the highest 5% Brysbaert et al. (2016).
2.3 Computational Use of Dictionaries
A lexicon with information about word meanings and usage is important to solve many NLP tasks. Doman-specific dictionaries for NLP have been developed over the years, initially through handcrafting and then automatic construction Riloff (1996); Lovis et al. (1998); Dolan et al. (1999); Przepiórkowski et al. (2014).
One interesting NLP problem tackled using dictionaries is word sense disambiguation (WSD), which explores the use of word sense in a particular context. Lexicographers study word senses in dictionaries Kilgarriff (1992) and suggest that computer readable dictionary corpora provide new opportunities for lexicography. Markowitz et al. Markowitz et al. (1986) and Montemagni et al. Montemagni and Vanderwende (1992) extract lexical information from implicit information in dictionary definitions. Dolan (1994) suggests that dictionary definitions are too fine-grained to use for WSD, and proposes a method to cluster word senses based on WordNet definitions and semantic relations. Pentheroudakis et al. (1994) identify morphologically related words in machine readable dictionaries (MRDs) and provide a certainty score for the extracted relations. Automatically gathered information from dictionaries have also been used to enrich thesauruses Oliveira and Gomes (2013).
2.4 Definitions and Word Embeddings
Word embeddings are a multi-dimensional vector representation of a word that conveys semantic and syntactic information. Words with similar meanings have embeddings closer to each other in the vector space. Popular word embeddings like Word2vec Mikolov et al. (2013), GloVe Pennington et al. (2014), and FastText Bojanowski et al. (2017) have been used for many NLP tasks.
Definition modeling seeks to generate word definitions given a word and its embedding Noraset et al. (2016). Definition generation solves the WSD problem by creating definitions instead of choosing a correct one Bevilacqua et al. (2021). Chang et al. analyze what contexualized word representations capture, and show these representations can be interpreted in a human-readable way Chang and Chen (2019). Huang et al. add a re-ranking mechanism to the decoder-encoder model to control specificity Huang et al. (2021). Liu et al use capsule networks first proposed in Sabour et al. (2017) to decompose word embeddings into morpheme-like vectors and reconstruct them into sense embeddings. Word embeddings have also been extracted from Urban dictionary Wilson et al. (2020a) and Twitter Wilson et al. (2020b).
3 Dataset
@@
Rate limitations on using the LLM APIs prevented us from studying the definitions of all English-language words, so we were selective in what words we study. Starting from a lexicon containing 246,591 English words sorted by their frequency, we sample 3000 words representing 500 words in each of six frequencies tiers:
  • Common: rank 1 to 500, and 501 to 1000.
  • Less Common: 5000-5500, and 10000-10500.
  • Rare: 20000-20500, and 50000-50500.
We cleaned the dataset by removing any word whose definition is missing in any source. This leaves 2512 words with complete definitions remaining in our final dataset. Table 1 presents a breakdown of these words by frequency tier and part of speech (POS) tag, demonstrating that our sampled words are well distributed by both features. Nouns are the most common type of words in the dataset.
Common Less common Rare Total
NOUN 749 475 182 1406
VERB 179 149 53 381
ADJ 153 77 19 249
ADV 83 32 16 131
Other 55 0 290 345
Total 1219 733 560 2512
Table 1: Word frequency and POS tag summary.
@@
  • Type 1: “What is the meaning of this word?"
  • Type 2: “Define this word."
This experimental design enables us to make a variety of comparisons between definitions produced (a) between different prompts on the same model, (b) two different models, namely GPT3.5 and GPT4.0, (c) comparisons between models and published dictionaries, (d) between common, moderate, and rare words, and (e) between different parts of speech.
To provide fair grounds for comparison, we extracted published definitions for each of these words from three online dictionaries, namely:
  • WordNet Miller (1995) (http://wordnetweb.princeton.edu/perl/webwn): This manually constructed database of word definitions and semantic relationships among words. It is highly accurate and serves as a useful tool in the field of NLP and linguistics.
  • Merriam-Webster dictionary Merriam-Webster Online (2023) (https://www.merriam-webster.com): The Merriam-Webster.com site is a commercial digital dictionary which is regularly updated.
  • Dictionary.com (https://www.dictionary.com): An online dictionary based on the Random House Unabridged Dictionary, and is updated regularly by editors.
The length distribution of these sources is summarized in Table 2, including mean and min/max lengths and standard deviation. Observe that GPT4 definitions are generally a little longer than those produced by GPT3.5, but within a given model the two prompts produce similar length texts. The correlation of the lengths of the two GPT3 definitions is 0.61, and that of the two GPT4 definitions is 0.74.
Two of the published dictionaries (Merriam-Webster and Dictionary.com) produce substantially longer definitions than WordNet and GPT3.5.
@@
Table 2: Length summary for collected dictionary datasets, showing the mean, min, max and standard deviation for each source, plus the number of queried lexicon terms defined within each sources (out of 3000).
4 Does ChatGPT Plagiarize?
Although neural network-based generative models synthesize response texts instead of explicitly cutting-and-pasting from training texts, this does not mean they cannot plagiarize, say through overfitting model parameters. Dictionary definitions provide an interesting domain to assess this possibility: they are precisely written texts that are readily available in training data. There is a long history of dictionaries plagiarizing each other Cooper (1962), and commonly include definitions of fake words (such as Mountweazel or jungftak) in order to catch wholesale pirating by other works Williams (2016).
Various plagiarism detection techniques Foltýnek et al. (2020); Zhou et al. (2022) can be used to check if GPT models are plagiarizing. The originality of essays produced by ChatGPT can be evaluated using plagiarism detection tools Khalil and Er (2023). The result indicates that ChatGPT output is complex enough to avoid being caught, and that ChatGPT itself performs better on determining whether the text is generated by the model or not.
We investigate this issue by identifying the longest common substring between the ChatGPT-generated definition for a given word and the corresponding definition in a published dictionary. Table 3 reports the words which contain the longest case-insensitive match between each model/dictionary pair. These examples are quite compelling, with common sequences as long as 17 words capturing the heart of the definition.
@@ 4 Does ChatGPT Plagiarize?
Table 3: The longest common substring between GPT-model generated definitions and published dictionaries. These matching phrases often capture the primary sense of the underlying word.
These examples are provocative, but not necessarily damning. Determining a meaningful expected longest common substring length in our instance is challenging, for several reasons. For randomly generated sequences, statisticians have proven this quantity grows logarithmically in the length of the sequences Arratia and Waterman (1985); Barros et al. (2019). But natural language text is far from random, and uses a large vocabulary as opposed to a constant-sized alphabet. Further, text with a high semantic similarity (two definitions of the same word) should share greater surface similarity than mismatched definitions. The length of the definitions matter, in a non-trivial way: longer, more detailed definitions should be expected to contain longer matches than briefer descriptions.
To assess whether these common phrases reflect untoward plagiarism or instead naturally constrained word choices in precise definitions, we compare the degree of borrowing between published dictionaries with what one sees generated by models. Standards for acceptable borrowing in published dictionaries are established by social convention and copyright law. Do LLMs borrow more from published dictionaries than dictionaries borrow from each other?
@@
Figure 1 presents the frequency distributions of match length between all three pairs of published dictionaries, and the two prompts for each model. They show that each pair of dictionaries share a small tail of long common substrings in their definitions, with substantially greater sharing between the two definitions from GPT-3 and GPT-4. There is nothing untoward about this, just that the generative models repeat language in the two equivalent definitions. Surprisingly, GPT-4 repeats itself substantially more frequently than GPT-3, perhaps an artifact from efforts to reduce model hallucination.
The average length of common strings, and frequencies of long matches reported in Table 4 shows more aggressive borrowing between Merriam-Webster and Dictionary.com than any other pairing of published dictionaries or dictionary-model pairs. In particular, the 155 long matches (of length $\geq 5$) observed between Merriam-Webster and Dictionary.com were roughly twice as many as between either of these dictionaries and any model. There seems no real evidence that GPT models unfairly replicate published training data, despite the long matches reported in Table 3.
Matched Pair
Average
longest
match
Count
matches
len5absent5$\geq$ 5≥ 5
Wordnet / Merriam-Webster 1.32 98
Wordnet / Dictionary.com 1.13 77
Merriam-Webster / Dictionary.com 1.54 155
GPT3-1 / Wordnet 1.55 99
GPT3-2 / Wordnet 1.59 97
GPT4-1 / Wordnet 1.65 93
GPT4-2 / Wordnet 1.62 91
GPT3-1 / Merriam-Webster 1.23 80
GPT3-2 / Merriam-Webster 1.25 75
GPT4-1 / Merriam-Webster 1.20 77
GPT4-2 / Merriam-Webster 1.18 70
GPT3-1 / Dictionary.com 1.00 69
GPT3-2 / Dictionary.com 1.00 71
GPT4-1 / Dictionary.com 1.04 65
GPT4-2 / Dictionary.com 1.04 75
GPT3-1 / GPT3-2 3.09 592
GPT4-1 / GPT4-2 4.00 711
Table 4: Average length and number of long matches ($\geq 5$ words) between definition sources.
5 Are GPT Definitions Consistent with Dictionaries?
@@
  • BERT Reimers and Gurevych (2019) – a modification of BERT that uses Siamese network structures to extract sentence-level representation using triplet loss.
  • MPnet Song et al. (2020) – a pre-trained model that combines permuted language modeling and the use of auxiliary position information to derive feature embedding.
  • Edit Distance – a string metric that represents the minimum number of insertions, deletions, or replacements required to transform one string into another. Here the edit distance is normalized by dividing itself by the maximum length of the two strings.
We calculate the distance between definitions provided by GPT models and online dictionaries. Table 5 shows the cosine distance between each pair of online dictionaries based on sentence embeddings and edit distance. The definitions from ChatGPT appear closer to Wordnet and Merriam-Webster, while definitions from GPT4 are closer to those on Dictionary.com.
Dictionary GPT model BERT MPNet
Edit
Dist.
Wordnet GPT3-1 0.33 0.43 4.72
GPT3-2 0.33 0.44 4.81
GPT4-1 0.37 0.49 5.01
GPT4-2 0.35 0.45 5.08
Merriam -Webster GPT3-1 0.31 0.38 4.49
GPT3-2 0.30 0.39 4.50
GPT4-1 0.33 0.43 4.53
GPT4-2 0.31 0.40 4.59
Dictionary .com GPT3-1 0.36 0.31 4.92
GPT3-2 0.35 0.31 4.86
GPT4-1 0.33 0.32 4.59
GPT4-2 0.31 0.28 4.63
Table 5: Distance between online dictionaries and definitions provided by GPT models (cosine).
Similarly, Table 6 reports the distance between each pair of GPT model definitions. For both GPT-3.5 and GPT-4, the definitions produces by the same model with different prompts are closest by all three distance metrics. Both BERT and MPNet-based distances perform well in capturing the similarity of definition pairs. We use cosine distance of BERT embeddings as our distance function in the rest of this paper, although similar results follow from MPNet.
Model 1 Model 2 BERT MPNet
Edit
Dist.
GPT3-1 GPT3-2 0.18 0.18 3.96
GPT4-1 0.28 0.24 4.79
GPT4-2 0.25 0.19 4.86
GPT3-2 GPT3-1 0.18 0.18 3.96
GPT4-1 0.26 0.22 4.61
GPT4-2 0.22 0.17 4.65
GPT4-1 GPT3-1 0.28 0.24 4.79
GPT3-2 0.26 0.22 4.61
GPT4-2 0.17 0.12 3.84
GPT4-2 GPT3-1 0.25 0.19 4.86
GPT3-2 0.22 0.17 4.65
GPT4-1 0.17 0.12 3.84
Table 6: Distance between definitions generated from GPT models (cosine). The two distinct definitions produced by the same version of GPT are closer than those generated by different models.
5.1 POS Analysis
A natural hypothesis is that certain classes of words are easier to generate reliable definitions for than others. Nouns represent objects, which might appear to be easier to precisely than descriptive words like adjectives and adverbs. We obtain a dominant POS tag on each word using the NLKT library Bird et al. (2009). Table 7 presents the average cosine distance error between generated and published definitions, for each model and POS type. Curiously, GPT3.5 proves most accurate on descriptive words, which GPT4 produces its best definitions for nouns. The results are consistent across all three published dictionaries, and indistinguishable for Type 1 and Type 2 prompts.
Model POS WordNet
Merriam-
Webster
Dictionary.
com
GPT3-1 NOUN 0.32 0.30 0.35
VERB 0.32 0.34 0.42
ADJ 0.33 0.28 0.35
ADV 0.29 0.28 0.41
GPT3-2 NOUN 0.33 0.30 0.34
VERB 0.32 0.33 0.40
ADJ 0.36 0.27 0.34
ADV 0.29 0.27 0.39
GPT4-1 NOUN 0.36 0.32 0.31
VERB 0.37 0.35 0.35
ADJ 0.40 0.33 0.33
ADV 0.45 0.40 0.37
GPT4-2 NOUN 0.35 0.30 0.30
VERB 0.35 0.33 0.34
ADJ 0.38 0.31 0.31
ADV 0.41 0.35 0.35
Table 7: Cosine distance based on BERT for different POS tags.
5.2 Word Frequency Analysis

### v3

@@ Abstract
Index Terms:
large language model, dictionary, word embedding
I Introduction
Many generations of readers have relied on the notion that the meaning of a word is what it says in the dictionary, but the primacy of dictionary definitions as the arbitrator of word meanings has come under threat by recent progress in natural language processing. Vector representations (word embeddings) have proven more valuable computationally than handcrafted semantics or definitions in describing the effective meaning of vocabulary words in a given language. And generative dialog systems like ChatGPT will happily produce full text definitions of every word when asked; indeed multiple versions of these definitions in response to different prompts. This paper address several distinct questions:
@@ I Introduction
  • LLMs have been described as “stochastic parrots” [1]. To what extent are the definitions they produce mere repetitions or even plagiarisms of classical dictionary definitions?
For over 2,500 carefully selected words, we extract definitions from three different dictionaries (WordNet, Merriam-Webster, and Random House/Dictionary.com), and also vector representations from two prominent collections of word embeddings (Glove[2] and FastText[3]). We also asked two versions of ChatGPT (3.5 and 4.0) for the definitions of every word, each with two different prompts. We measure the distances between these representations using techniques including vector distance/similarity, edit distance, and neighborhood correlations.
@@
  • The surface form of generated definitions differ substantially from those of published dictionaries – Although ChatGPT presumably trained in part on the handcrafted definitions in our study, little-to-no unexpected trace of these texts remain in the generated text.
@@
  • ChatGPT generated word definitions are highly accurate, consistent with published dictionaries – Of the 50 most distant pairs between Merriam-Webster and GPT4 definitions based on their definition embeddings, human evaluation confirms that GPT4 recognized 48 compared to 34 for GPT3, and that all but one of these definitions matched the primary sense of the published dictionary.
@@
  • The consistency of LLM-generated definitions is relatively independent of frequency, unlike word embeddings – We propose a new average distance correlation metric to compare word embeddings in different dimensional spaces. Our results suggest that SBERT embeddings of definitions (synthesized or published) may create more accurate word embeddings for low frequency words than traditional methods.
@@
This paper is organized as follows. Section II presents related work on LLMs and word embeddings. The dictionaries and generative models we use to formulate our dataset are described in Section III. The question of just how originial generated definitions are is considered in Section IV. We use vector representations to quantify the similarity of definitions by source in Section V, and between definitions and word embeddings in Section VI. We conclude with directions for future work in Section VII.
II Related Work
II-A Large Language Models
@@
Large language models (LLM) recently gained popularity, with ChatGPT reaching 100 million users in only three months [4]. Generative Pre-trained Transformer (GPT) models are a series of large language models developed by OpenAI. Each model in the series [5, 6, 7, 8] is trained on a larger corpus of text and achieves better results on natural language processing (NLP) tasks compared to its predecessors. LLMs have had significant impact across many domains, including medical imaging diagnostics [9], recommendation systems [10], and education [11]. They have been effectively applied to a variety of tasks such as taxonomy classification [12], translation quality assessment [13], and cloud configuration generation [14].
II-B Word Embeddings
Text embeddings provide a powerful representation of text data, and have been widely used across various applications, including bot detection [15], mortality prediction [16, 17], and other significant tasks. Word embeddings, in particular, represent words as multi dimensional vectors that encode both semantic and syntactic information. Words with similar meanings are positioned closer together in the vector space, making these embeddings highly effective for a range of NLP tasks. Prominent examples of word embeddings include Word2Vec [18], GloVe [2], and FastText [3], which have been widely adopted in the NLP community.
III Dataset
@@
Rate limitations on using the LLM APIs prevented us from studying the definitions of all English-language words, so we were selective in what words we study. Starting from a lexicon containing 246,591 English words sorted by their frequency, we sampled 3000 words by including the 1000 most frequent words and additional words drawn from four frequency tiers: 5000-5500, 10000-10500, 20000-20500, and 50000-50500. We cleaned the dataset by removing any word whose definition is missing in any source. This leaves 2512 words with complete definitions remaining in our final dataset.
@@
  • Type 1: “What is the meaning of this word?”
  • Type 2: “Define this word.”
This experimental design enables us to make a variety of comparisons between definitions produced (a) between different prompts on the same model, (b) two different models, namely GPT3.5 and GPT4.0, (c) comparisons between models and published dictionaries, (d) between common, moderate, and rare words, and (e) between different parts of speech. All the experiments in this paper are performed on a server with Intel Xeon Gold 6140 CPU (2.30GHz, 36 cores, 25MB Cache).
To provide fair grounds for comparison, we extracted published definitions for each of these words from three online dictionaries, namely: WordNet [19], Merriam-Webster dictionary  [^1] https://www.merriam-webster.com/, and Dictionary.com  [^2] https://www.dictionary.com.
The definition length distribution of these sources is summarized in Table I, including mean and min/max lengths and standard deviation. Observe that GPT4 definitions are generally a little longer than those produced by GPT3.5, but within a given model the two prompts produce similar length texts. The correlation of the lengths of the two GPT3 definitions is 0.61, and that of the two GPT4 definitions is 0.74. Two of the published dictionaries (Merriam-Webster and Dictionary.com) produce substantially longer definitions than WordNet and GPT3.5.
@@
TABLE I: Definition length summary for collected dictionary datasets, showing the mean, min, max and standard deviation for each source, plus the number of queried lexicon terms defined within each sources (out of 3000).
IV Does ChatGPT Plagiarize?
Although neural network-based generative models synthesize response texts instead of explicitly cutting-and-pasting from training texts, this does not mean they cannot plagiarize, say through overfitting model parameters. Dictionary definitions provide an interesting domain to assess this possibility: they are precisely written texts that are readily available in training data. We investigate this issue by identifying the longest common substring between the ChatGPT-generated definition for a given word and the corresponding definition in a published dictionary. Table II reports the words which contain the longest case-insensitive match between each model/dictionary pair. These examples are quite compelling, with common sequences as long as 17 words capturing the heart of the definition.
@@ IV Does ChatGPT Plagiarize?
TABLE II: The longest common substring between GPT-model generated definitions and published dictionaries. These matching phrases often capture the primary sense of the underlying word.
Determining a meaningful expected longest common substring length in our instance is challenging, for several reasons. For randomly generated sequences, statisticians have proven this quantity grows logarithmically in the length of the sequences [20, 21]. But natural language text is far from random, and uses a large vocabulary as opposed to a constant-sized alphabet. Further, text with a high semantic similarity (two definitions of the same word) should share greater surface similarity than mismatched definitions. The length of the definitions matter, in a non-trivial way: longer, more detailed definitions should be expected to contain longer matches than briefer descriptions.
To assess whether these common phrases reflect untoward plagiarism or instead naturally constrained word choices in precise definitions, we compare the degree of borrowing between published dictionaries with what one sees generated by models. Standards for acceptable borrowing in published dictionaries are established by social conventions and copyright law. Figure 1 presents the frequency distributions of match length between all three pairs of published dictionaries, and the two prompts for each model. They show that each pair of dictionaries share a small tail of long common substrings in their definitions, with substantially greater sharing between the two definitions from GPT-3 and GPT-4. There is nothing untoward about this, just that the generative models repeat language in the two equivalent definitions. Surprisingly, GPT-4 repeats itself substantially more frequently than GPT-3, perhaps an artifact from efforts to reduce model hallucination. The average length of common strings, and frequencies of long matches shows more aggressive borrowing between Merriam-Webster and Dictionary.com than any other pairing of published dictionaries or dictionary-model pairs. In particular, 155 long matches of length $\geq 5$ are observed between Merriam-Webster and Dictionary.com, which is roughly twice as many as between either of these dictionaries and any model. There seems no real evidence that GPT models unfairly replicate published training data, despite the long matches reported in Table II.
@@
V Are GPT Definitions Consistent with Dictionaries?
@@
  • SBERT [22] – a modification of BERT that uses Siamese network structures to extract sentence-level representation using triplet loss.
  • MPnet [23] – a pre-trained model that combines permuted language modeling and the use of auxiliary position information to derive feature embedding.
  • Edit Distance – a string metric that represents the minimum number of insertions, deletions, or replacements required to transform one string into another. Here the edit distance is normalized by dividing the maximum length of the two strings.
We calculate the distance between definitions provided by GPT models and online dictionaries. The definitions from ChatGPT appear closer to Wordnet and Merriam-Webster, while definitions from GPT4 are closer to those on Dictionary.com. We use cosine distance of SBERT embeddings as our distance function in the rest of this paper, although similar results follow from MPNet.
V-A POS Analysis
A natural hypothesis is that certain classes of words are easier to generate reliable definitions for than others. Nouns represent objects, which might appear to be easier to precisely than descriptive words like adjectives and adverbs. We obtain a dominant POS tag on each word using the NLTK library [24]. The average cosine distance error between generated and published definitions is calculated for each model and POS type. Curiously, GPT3.5 proves most accurate on descriptive words, which GPT4 produces its best definitions for nouns. The results are consistent across all three published dictionaries, and indistinguishable for Type 1 and Type 2 prompts.
V-B Word Frequency Analysis